### 25 Mar 2026 v1

@@ 1.2. Finiteness properties
The use of brackets versus parentheses is inspired by the notation of Bader–Sauer for higher Kazhdan properties [BS]. Note that the condition in the second item is trivially satisfied for any $T\subseteq S$ with $|T|\geq n$. Actions of type $\operatorname{A}_{n}]$ have an established tight connection to finiteness properties, thanks to permutational wreath products. Indeed, for an infinite group $B$ of type $\operatorname{\mathcal{F}}_{n}$, the permutational wreath product $B\wr_{S}G$ is of type $\operatorname{\mathcal{F}}_{n}$ if and only if the action $G\curvearrowright S$ is of type $\operatorname{A}_{n}] [$dC06, BdCK15, FFKLZ].
@@ 1.3. Embeddings
Here the “hence” in the parenthetical follows from Theorems A and B and Lemma 5.2. It would be interesting to know whether the converse of this parenthetical is true, i.e., whether every finitely presented abstract twisted Brin–Thompson group sharply embeds in a normal pair of the form $(G,\ker(G\curvearrowright S))$ for some $G\curvearrowright S$ of type $\operatorname{A}_{2}]$ (see Question 5.8).
As in the case of the classical Boone–Higman conjecture, these conjectures would also provide an elegant characterization of solvability of the word problem, thanks to essentially the same algorithm: see Proposition 5.4. In particular the converses of these conjectures are true. We should emphasize that the desired embeddings in these relative conjectures need to be sharp. Indeed, every finitely presented group $G$ (regardless of the word problem) embeds non-sharply in a group admitting a type $\operatorname{A}_{2}]$ action, namely $G$ acting trivially on a point.
@@ 1.4. Other properties
This paper is organized as follows. In Section 2 we define abstract twisted Brin–Thompson groups. In Section 3 we prove Theorem A, that they are relatively simple. In Section 4 we prove Theorem B regarding finiteness properties. In Section 5 we investigate embedding properties and prove Theorem C. In Section 6 we investigate geometric properties, including actions on hyperbolic spaces. In Section 7 we look at homological properties, and prove that every abstract twisted Brin–Thompson group is boundedly acyclic and $\ell^{2}$-invisible. Finally, in Section 8 we characterize precisely which abstract twisted Brin–Thompson groups are $C^{*}$-simple.
@@ Acknowledgments
This work has benefited from discussions with Jim Belk, Collin Bleak, Matt Brin, Kai-Uwe Bux, James Hyde, Adrien Le Boudec, Francesco Matucci, Roman Sauer, and Jakub Tucker. FFF is supported by the Herchel Smith Postdoctoral Fellowship Fund.
@@ 5. Embedding results
The action $G\curvearrowright S$ constructed in the proof above is of type $\operatorname{A}_{2}]$ but not of type $\operatorname{A}_{\infty}]$ in general, for example when $\Gamma$ is a (finitely presented, simple) Burger–Mozes group [FFKLZ, Corollary 1.6].
We should emphasize that one advantage of allowing non-faithful actions is that, a priori, the simple quotient obtained by taking $SV_{G}$ modulo its largest subgroup is not necessarily finitely presented, and so even though sharpness of the embedding induces an embedding into the quotient, this quotient is not necessarily useful for Boone–Higman-related purposes. (That being said, we do not know an explicit example where it fails, see Question 5.9.) This is in contrast to other examples of relatively simple groups, for instance in the braided Thompson group or labeled Thompson groups the quotient by the largest subgroup is $V$, which is finitely presented. Also, for the relatively simple lamplighters in Lemma 3.1, if the wreath product is finitely presented then so is the quotient by the largest normal subgroup, since the quotient splits.
It is not difficult to see that the converse of the (relative) (permutational) Boone–Higman conjecture is true, as the following shows.
Proposition 5.4.
Let $\Gamma$ be a finitely generated group that sharply embeds in a finitely presented relatively simple group. Then $\Gamma$ has solvable word problem.
Proof.
Let $(G,N)$ be a finitely presented relatively simple normal pair such that $(\Gamma,\{1\})$ embeds in $(G,N)$, and we have to prove $\Gamma$ has solvable word problem.
We follow the classical Kuznetsov algorithm [Kuz58], which treated the case $N=\{1\}$. Fix a finite presentation $\langle S\mid R\rangle$ for $G$, identify $\Gamma$ with a subgroup of $G$, and fix a word $w$ in the alphabet $S$ representing an element of $\Gamma$. We run a “day and night” algorithm, which returns “yes” if $w$ represents the identity, and “no” otherwise.
We claim that this algorithm terminates. Indeed, if $w$ represents the identity, then it must eventually appear in an enumeration of all the relations. Otherwise, $w$ represents an element in $\Gamma\setminus\{1\}\subset G\setminus N$. By relative simplicity, since $\langle\hskip-1.99168pt\langle w\rangle\hskip-1.99168pt\rangle$ is not contained in $N$ it must equal all of $G$, so $\langle S\mid R\cup\{w\}\rangle$ is a presentation for the trivial group. Thus, all the generators in $S$ must eventually appear in an enumeration of all the relations of this new presentation. ∎
Note that a finitely presented relatively simple group itself may have unsolvable word problem, for example if $|S|=1$ and $G$ has unsolvable word problem, then $SV_{G}$ has unsolvable word problem (but any group sharply embedding in it has solvable word problem). As a remark, the braided Thompson group $bV$ is an example of a finitely presented relatively simple group with solvable word problem for which the Boone–Higman conjecture (and even the relative Boone–Higman conjecture) remains open.
@@
First assume (i) holds, so $G\curvearrowright S$ is of type $\operatorname{A}_{2}]$ and $\Gamma$ sharply embeds in $(G,\ker(G\curvearrowright S))$, say via $\iota\colon\Gamma\to G$. Note that by Observation 5.5, we could assume the action is transitive by just restricting to a single orbit, but then the embedding of $\Gamma$ might no longer be sharp, so more work is needed. Let $S_{1},\dots,S_{n}\subseteq S$ be the $G$-orbits of $G\curvearrowright S$. Consider the action of the finitely presented group $G^{n}$ on $S_{1}\times\cdots\times S_{n}$ via $(g_{1},\dots,g_{n}).(s_{1},\dots,s_{n})\coloneqq(g_{1}.s_{1},\dots,g_{n}.s_{n})$, so this action is transitive. Since $G\curvearrowright S$ has finitely many orbits of pairs, the same is true of this action of $G^{n}$ on $S_{1}\times\cdots\times S_{n}$. Since point stabilizers for $G\curvearrowright S$ are finitely generated, the same is true for the new action. Finally, embed $\Gamma$ into $G^{n}$ diagonally via $\iota$ in each factor, and notice that since each non-trivial element of $\Gamma$ must act non-trivially on some $S_{i}$, the embedding is sharp relative the kernel of the action of $G^{n}$. This shows (ii).
@@
Note that a “yes” answer to the first part of Question 5.10 would tell us that (relPBH) implies (PBH), and the second part would tell us that (relBH) implies (BH). By Theorem C the first part would imply that (BH), (relPBH), and (PBH) are all equivalent to each other, and the second part would imply that (BH), (relPBH), and (relBH) are all equivalent to each other.

### 02 Apr 2026 v2

@@ 1.2. Finiteness properties
The use of brackets versus parentheses is inspired by the notation of Bader–Sauer for higher Kazhdan properties [BS]. Note that the condition in the second item is trivially satisfied for any $T\subseteq S$ with $|T|\geq n$. Actions of type $\operatorname{A}_{n}]$ have an established tight connection to finiteness properties, thanks to permutational wreath products. Indeed, for an infinite group $B$ of type $\operatorname{\mathcal{F}}_{n}$, the permutational wreath product $B\wr_{S}G$ is of type $\operatorname{\mathcal{F}}_{n}$ if and only if the action $G\curvearrowright S$ is of type $\operatorname{A}_{n}] [$dC06, BdCK15, FFKLZ25].
@@ 1.3. Embeddings
Here the “hence” in the parenthetical follows from Theorems A and B and Lemma 5.2. It would be interesting to know whether the converse of this parenthetical is true, i.e., whether every finitely presented abstract twisted Brin–Thompson group sharply embeds in a normal pair of the form $(G,\ker(G\curvearrowright S))$ for some $G\curvearrowright S$ of type $\operatorname{A}_{2}]$ (see Question 5.7).
As in the case of the classical Boone–Higman conjecture, these conjectures would also provide an elegant characterization of solvability of the word problem, thanks to essentially the same algorithm; in fact the relative Boone–Higman conjecture, the weakest of the four, is a natural strengthening of a characterization of groups with solvable word problem that follows from the Boone–Higman–Thompson theorem, and that we present in Appendix A (see Proposition A.4). In particular the converses of these conjectures are true. We should emphasize that the desired embeddings in these relative conjectures need to be sharp. Indeed, every finitely presented group $G$ (regardless of the word problem) embeds non-sharply in a group admitting a type $\operatorname{A}_{2}]$ action, namely $G$ acting trivially on a point.
@@ 1.4. Other properties
This paper is organized as follows. In Section 2 we define abstract twisted Brin–Thompson groups. In Section 3 we prove Theorem A, that they are relatively simple. In Section 4 we prove Theorem B regarding finiteness properties. In Section 5 we investigate embedding properties and prove Theorem C. In Section 6 we investigate geometric properties, including actions on hyperbolic spaces. In Section 7 we look at homological properties, and prove that every abstract twisted Brin–Thompson group is boundedly acyclic and $\ell^{2}$-invisible. Finally, in Section 8 we characterize precisely which abstract twisted Brin–Thompson groups are $C^{*}$-simple. In Appendix A we prove a strengthening of the Boone–Higman–Thompson theorem (Proposition A.4) that has an interesting connection to the relative Boone–Higman conjecture.
@@ Acknowledgments
This work has benefited from discussions with Carolyn Abbott, Jim Belk, Collin Bleak, Matt Brin, Kai-Uwe Bux, James Hyde, Adrien Le Boudec, Robbie Lyman, Francesco Matucci, Roman Sauer, and Jakub Tucker. FFF is supported by the Herchel Smith Postdoctoral Fellowship Fund.
@@ 5. Embedding results
The action $G\curvearrowright S$ constructed in the proof above is of type $\operatorname{A}_{2}]$ but not of type $\operatorname{A}_{\infty}]$ in general, for example when $\Gamma$ is a (finitely presented, simple) Burger–Mozes group [FFKLZ25, Corollary 1.6].
We should emphasize that one advantage of allowing non-faithful actions is that, a priori, the simple quotient obtained by taking $SV_{G}$ modulo its largest subgroup is not necessarily finitely presented, and so even though sharpness of the embedding induces an embedding into the quotient, this quotient is not necessarily useful for Boone–Higman-related purposes. (That being said, we do not know an explicit example where it fails, see Question 5.8.) This is in contrast to other examples of relatively simple groups, for instance in the braided Thompson group or labeled Thompson groups the quotient by the largest subgroup is $V$, which is finitely presented. Also, for the relatively simple lamplighters in Lemma 3.1, if the wreath product is finitely presented then so is the quotient by the largest normal subgroup, since the quotient splits.
It is not difficult to see that the converse of the (relative) (permutational) Boone–Higman conjecture is true, in fact the relative Boone–Higman conjecture is a natural strengthening of a characterization of groups with solvable word problem that follows from the Boone–Higman–Thompson theorem, and that we present in the appendix (Proposition A.4). Note that a finitely presented relatively simple group itself may have unsolvable word problem, for example if $|S|=1$ and $G$ has unsolvable word problem, then $SV_{G}$ has unsolvable word problem (but any group sharply embedding in it has solvable word problem). As a remark, the braided Thompson group $bV$ is an example of a finitely presented relatively simple group with solvable word problem for which the Boone–Higman conjecture (and even the relative Boone–Higman conjecture) remains open.
@@
First assume (i) holds, so $G\curvearrowright S$ is of type $\operatorname{A}_{2}]$ and $\Gamma$ sharply embeds in $(G,\ker(G\curvearrowright S))$, say via $\iota\colon\Gamma\to G$. Note that by Observation 5.4, we could assume the action is transitive by just restricting to a single orbit, but then the embedding of $\Gamma$ might no longer be sharp, so more work is needed. Let $S_{1},\dots,S_{n}\subseteq S$ be the $G$-orbits of $G\curvearrowright S$. Consider the action of the finitely presented group $G^{n}$ on $S_{1}\times\cdots\times S_{n}$ via $(g_{1},\dots,g_{n}).(s_{1},\dots,s_{n})\coloneqq(g_{1}.s_{1},\dots,g_{n}.s_{n})$, so this action is transitive. Since $G\curvearrowright S$ has finitely many orbits of pairs, the same is true of this action of $G^{n}$ on $S_{1}\times\cdots\times S_{n}$. Since point stabilizers for $G\curvearrowright S$ are finitely generated, the same is true for the new action. Finally, embed $\Gamma$ into $G^{n}$ diagonally via $\iota$ in each factor, and notice that since each non-trivial element of $\Gamma$ must act non-trivially on some $S_{i}$, the embedding is sharp relative the kernel of the action of $G^{n}$. This shows (ii).
@@
Note that a “yes” answer to the first part of Question 5.9 would tell us that (relPBH) implies (PBH), and the second part would tell us that (relBH) implies (BH). By Theorem C the first part would imply that (BH), (relPBH), and (PBH) are all equivalent to each other, and the second part would imply that (BH), (relPBH), and (relBH) are all equivalent to each other.
@@ 8. $C^{*}$-simplicity
Appendix A A strengthening of the Boone–Higman–Thompson theorem
Recall from Remark 5.6 the Boone–Higman–Thompson theorem, that a finitely generated group has solvable word problem if and only if it embeds in a finitely generated simple group that embeds in a finitely presented group [BH74, Tho80]. In this appendix we slightly strengthen this theorem, and also discuss connections to twisted Brin–Thompson groups and the Boone–Higman conjecture. The statement can be phrased in a number of equivalent ways: the following is left to the reader.
Lemma A.1.
Let $\Gamma\leq P$ be groups. Then the following are equivalent.
  • Every non-identity element of $\Gamma$ normally generates $P$.
  • Every proper normal subgroup of $P$ intersects $\Gamma$ trivially.
  • Every non-trivial quotient of $P$ is injective on $\Gamma$. ∎
We will need the following lemma.
Lemma A.2.
Let $\Gamma$ be a finitely presented group and $1\neq g\in\Gamma$. Then there exists a finitely presented group $P$ containing $\Gamma$, such that $g$ normally generates $P$.
Proof.
First embed $\Gamma$ into a finitely generated simple group $G$, for instance the twisted Brin–Thompson group $\Gamma V_{\Gamma}$. We claim that there exists a finitely presented cover $P\to G$ satisfying the hypotheses. Indeed, since $\Gamma$ is finitely presented, only finitely many relations from $G$ suffice to ensure that $\Gamma$ embeds. Moreover, $g$ normally generates $G$ since $G$ is simple, that is, every generator of $G$ is a product of conjugates of powers of $g$, and this can again be expressed by only finitely many relations. ∎
Remark A.3.
Recently, Chatterji–Kassabov proved that $P$ can moreover be chosen to have property (T) [CK], hence property (T) can be assumed of the container group in Proposition A.4. In fact, via an easy construction using small cancellation theory over acylindrically hyperbolic groups [Hul16], the container group $P$ may be chosen to be a quotient of any given finitely presented acylindrically hyperbolic group. Since these additional properties are not relevant to us, we only need Lemma A.2, for which we could present a self-contained elementary proof.
The main result of this appendix is the following.
Proposition A.4.
Let $\Gamma$ be a finitely generated group. Then $\Gamma$ has solvable word problem if and only if there exists a finitely presented group $P$ such that $\Gamma$ embeds as a subgroup of $P$ trivially intersecting every proper normal subgroup of $P$.
Proof.
First suppose $\Gamma$ has solvable word problem, so by the Boone–Higman–Thompson theorem there exist a finitely generated simple group $G$ and a finitely presented group $P$ such that $\Gamma\leq G\leq P$. By Lemma A.2, there exists a finitely presented group $Q$ containing $P$ such that an element $1\neq g\in G$ normally generates $Q$. Since $G$ is simple, it follows that every non-identity element in $G$, hence in particular in $\Gamma$, normally generates $Q$. We conclude by Lemma A.1.
Now suppose $\Gamma$ is a subgroup of a finitely presented group $P$ such that $\Gamma$ trivially intersects every non-trivial normal subgroup of $P$. We must show that $\Gamma$ has solvable word problem.
We follow the classical Kuznetsov algorithm [Kuz58], which treated the case when $P$ is simple. Fix a finite presentation $\langle S\mid R\rangle$ for $P$, and fix a word $w$ in the alphabet $S$ representing an element of $\Gamma$. We run a “day and night” algorithm, which returns “yes” if $w$ represents the identity, and “no” otherwise.
We claim that this algorithm terminates. Indeed, if $w$ represents the identity, then it must eventually appear in an enumeration of all the relations. Otherwise, $w$ represents an element in $\Gamma\setminus\{1\}$, hence a normal generator of $P$, by Lemma A.1. Therefore $\langle S\mid R\cup\{w\}\rangle$ is a presentation for the trivial group. Thus, all the generators in $S$ must eventually appear in an enumeration of all the relations of this new presentation, and we are done. ∎
This immediately gives us the following strengthening of the Boone–Higman–Thompson theorem:
Corollary A.5.
A finitely generated group $\Gamma$ has solvable word problem if and only if it embeds in a finitely generated simple group $G$ that embeds in a finitely presented group $P$, such that $G$ normally generates PP.∎
The relative Boone–Higman conjecture amounts to saying that the group $P$ in Proposition A.4 can always be taken to be relatively simple, so conjecturally the statement about trivially intersecting every proper normal subgroup boils down to just trivially intersecting one proper normal subgroup (the largest). Note that if $\Gamma$ is infinite, then since it embeds in every non-trivial quotient of $P$ by Lemma A.1, it follows that $P$ has no non-trivial finite quotients. Thus a Zorn’s lemma argument going back to Higman [Hig51] shows that $P$ admits a (finitely generated) simple quotient. If one could show that, in this situation, $P$ admits a finitely presented simple quotient, then this would prove the Boone–Higman conjecture.